During a fireworks display, a shell is shot into air at an initial speed of 70 m/s at Angie of 75above the horizontal. The shell ignites once it reaches its highest point. How long did it take to ignite from the time it was launched? answer in seconds
1
Expert's answer
2020-07-06T15:15:10-0400

According to the


https://books.google.com.ua/books?id=QwzwfMr8qncC&pg=PA42&lpg=PA42&dq=a+body+is+thrown+at+some+angle+from+the+ground&source=bl&ots=R_M9bVR11W&sig=ACfU3U2YLUIzaQmA6YBa0BcaKK-Pc4UjgA&hl=ru&sa=X&ved=2ahUKEwjzmtP-p7bqAhVn2aYKHT2DBwYQ6AEwC3oECAkQAQ#v=onepage&q=a%20body%20is%20thrown%20at%20some%20angle%20from%20the%20ground&f=false


the x-coordinate of the shell changes by the following law:


"x = v_0t\cos\theta"

where "v_0 = 70m\/s" is the initial velocity, "t" is time and "\theta = 75\degree" is the launch angle.

The maximum distance that the shell can cover in x-direction is:


"x_M = \dfrac{v_0^2}{g}\sin (2\theta)"

Due to the symmetry of the problem, the shell reaches the maximum height when it reaches exactly half of the "x_M". Thus, substituting the "x = 0.5x_M" into the first equation and expressing "t", obtain


"\dfrac{v_0^2}{2g}\sin (2\theta) = v_0t\cos\theta\\
t = \dfrac{v_0\sin (2\theta) }{2g\cos\theta} = \dfrac{v_02\sin \theta\cos\theta }{2g\cos\theta} = \dfrac{v_0\sin \theta }{g}"

Substituting numerical values, get:


"t = \dfrac{70\cdot \sin 75\degree }{9.8} \approx 6.9s"

Answer. 6.9 s.